Nancy has built a very profitable business, and she credits the entrepreneurship program at her alma...

  1. Nancy has built a very profitable business, and she credits the entrepreneurship program at her alma mater for a lot of her success. She would like to donate money to her old school to help one worthy graduate each year establish his or her own business. She will donate the money today, with the understanding that the first award will go to a graduate of this year's freshman class. (That is, the first award will be made four years from now.) Her alma mater is able to invest the funds at a constant, annual, tax-free rate of 12%. How much must her donation be if she would like for the annual award to be $10,000 a year and wants the program to continue forever, even after she is no longer around?

    1. $52,960

    2. $59,315

    3. $83,333

    4. $93,333

    5. None of the above.

answer is B, but I don't get why the nper=3. I think it should be 4?
Answer #1
Amount required at t3 = 10000/12% = $            83,333
Amount of donation to be made now = 83333.33/1.12^3 = $            59,315